On any given days, the chance of you waking up a feel awesome is about 55%. You decide to keep track of your mood for 2 weeks (14 days): What is the chance that you will not feel awesome for exactly 10 days? (Write your answer with 4 decimal places) and What is the chance of you will feel awesome between 7 - 9 days. Round answer to 4 decimal places
Answer #1

This will be a binomial distribution with parameters:

n = 14, p = P(Awesome feel) = 0.55

q = 1 - p = 1 - 0.55 = 0.45

So,

a) P(Not feel awesome for exactly 10 days)

= P(Feel awesome for exactly 4 days)

= binom.dist(4, 14, 0.55, False) [Excel Formula]

= 0.0312

b) P(Between 7 and 9 days)

= binom.dist(9, 14, 0.55, True) - binom.dist(6, 14, 0.55, True) [Excel Formula]

= 0.5742